OpenStudy (anonymous):

add simplify by collecting like radical terms if possible: 8√80 + 2√45 =38 √5 is my answer

jimthompson5910 (jim_thompson5910):

Good job, you got the right answer.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

k

OpenStudy (anonymous):

it may be right but it seems large. let me check

jimthompson5910 (jim_thompson5910):

It's correct. I verified the answer myself.

OpenStudy (anonymous):

oh never mind, i am sorry.

OpenStudy (anonymous):

ok sorry. you got it.

OpenStudy (anonymous):

i loved pop 1280 btw

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

this is a site to help me also

OpenStudy (anonymous):

satellite73 need to double check my answer pls: subtract and simplify collect like terms if possible 9^√6 - 7^3√6 = 2^3√6 is my ans. is this correct

OpenStudy (anonymous):

satellite73 need to double check my answer pls: subtract and simplify collect like terms if possible 9^√6 - 7^3√6 = 2^3√6 is my ans. is this correct

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\[\sqrt[9]{6}-7\sqrt[3]{6}\] ? these are not like terms so you cannot subtract them

jimthompson5910 (jim_thompson5910):

Is it \[9\sqrt[3]{6}-7\sqrt[3]{6}\]??

jimthompson5910 (jim_thompson5910):

If so, then jazil you have the correct answer.

OpenStudy (anonymous):

yes thanks jim

OpenStudy (anonymous):

sorry satellite i left out the 9^3√6 - 7^3√6